# Stability and List-Replicability for Agnostic Learners

### v1

@@ 1 Introduction
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\bm{S}))\leq\mathcal{L}_{\mathcal{D}}(\mathcal{H})+\epsilon]\geq 1-\delta\ \text{where }n=n(\epsilon,\delta).$
@@
Stability.
Replicability is a fundamental principle of the scientific method. A study is replicable if it consistently yields the same results when repeated with new data drawn from the same distribution or source. In recent years, machine learning has seen a growing need to address the replication crisis [Bal23, Bak16]. Impagliazzo, Lei, Pitassi, and Sorrel [ILPS22] initiated a formal theoretical framework for studying replicability in machine learning. Since their work, a rapidly growing body of research has emerged that introduced various notions of replicability. These works and subsequent research showed that many of these notions of replicability are essentially equivalent. Furthermore, they established deep connections to other foundational concepts in learning theory, such as differential privacy [CMY23, BGH+23, KKMV23, GKM21a, GKM21b, CCMY24]. Additionally, a growing body of work has explored replicability in many data analysis and learning settings [ILPS22, BGH+23, KVYZ23, EKK+23, EKM+23, EHKS24, KKL+24, KKMV23].
In this paper, we focus on the notion of replicability where the learning algorithm is expected to often produce the same predictor when applied to two independent and identically distributed inputs. This concept was first introduced under the term global stability in [BLM20] and has since been refined and explored in subsequent works [GKM21a, KKMV23, CMY23, CCMY24]. We start by defining stability and global stability.
Definition 1.1 (Stability).
Given a function $\rho:(0,1)\rightarrow(0,1)$, a learning rule $\bm{\mathcal{A}}$ is a $\rho$-stable learner for a hypothesis class $\mathcal{H}$ if the following holds. For every $\epsilon>0$, there exists $n=n(\epsilon)$ such that for every realizable distribution $\mathcal{D}$, there exists a hypothesis $h$ satisfying
@@
Similarly, we call $\bm{\mathcal{A}}$ a $\rho$-stable agnostic learner for $\mathcal{H}$, if there exists $n=n(\epsilon)$ such that for every distribution $\mathcal{D}$ on $X\times\{0,1\}$, there exists a hypothesis $h\in\{0,1\}^{X}$ that satisfies (1) and
@@
To simplify terminology, we use the term $\rho$-stable to describe a hypothesis class $\mathcal{H}$ with a $\rho$-stable learner. Likewise, we call $\mathcal{H}$ agnostically $\rho$-stable if it has a $\rho$-stable agnostic learner.
Definition 1.2 (Global Stability).
We say that a hypothesis class $\mathcal{H}$ is globally stable if it is $\rho$-stable for a fixed constant $\rho\in(0,1)$. Similarly, a hypothesis class $\mathcal{H}$ is agnostically globally stable if it is agnostically $\rho$-stable for such a constant.
@@
Bun, Livni, and Moran [BLM20] showed that in the realizable setting, global stability is fully characterized by bounded Littlestone dimension (Definition 1.12). This result, combined with the seminal works of [Lit88, ABL+22], shows that global stability is equivalent to online learnability and approximately private learnability, as well as, some other notions of replicability [KKMV23, GKM21a, BGH+23].
@@
This striking result shows that agnostic global stability is far more restrictive than its realizable counterpart. Since finite classes are trivially stable, Theorem 1.3 shows that agnostic global stability is too restrictive to lead to interesting learnability phenomena. To remedy this, Chase et al. [CCMY24] introduced two relaxations of agnostic global stability and proposed a study of which hypothesis classes can be learned under these relaxed notions of stability.
@@
The first suggested relaxation, coincides with our definition of $\rho$-stability in Definition 1.1. A hypothesis class $\mathcal{H}$ is called excess-error dependent stable if it is agnostically $\rho$-stable for some $\rho:(0,1)\to(0,1)$. Here the excess-error refers to the parameter $\epsilon$ in Eq. 2.
Our main theorem provides a complete characterization of such classes. We show that a hypothesis class is $\rho$-agnostically stable learnable for some $\rho$ if and only if it has a bounded Littlestone dimension. We denote the Littlestone dimension of $\mathcal{H}$ as $\operatorname{Ldim}(\mathcal{H})$.
@@
    If $\operatorname{Ldim}(\mathcal{H})=\infty$, then $\mathcal{H}$ is not $\rho$-stable for any $\rho:(0,1)\to(0,1)$.
@@
    If $\operatorname{Ldim}(\mathcal{H})<\infty$, then $\mathcal{H}$ is agnostically $\rho$-stable for some $\rho:(0,1)\to(0,1)$.
Note that Theorem 1.4 (i) states that if $\operatorname{Ldim}(\mathcal{H})=\infty$, then even in the realizable case, we cannot achieve $\rho$-stability for any $\rho:(0,1)\to(0,1)$. This strengthens the result of Bun, Livni, and Moran [BLM20], which only overrules $\rho$-stability when $\rho>0$ is a fixed constant.
Combined with the work of [ABL+22], Theorem 1.4 implies that agnostic $\rho$-stability is equivalent to global stability, as well as to approximate private learnability and online learnability.
@@
Definition 1.5 (Class-error Dependent Stability).
@@
In learning theory, stability is more useful when paired with a guarantee that the learner typically outputs a hypothesis with a low population loss. However, the definition of global stability (Definition 1.2) only requires that the learner outputs a low-error hypothesis with some probability $\rho>0$, allowing the learner to output hypotheses with large population loss with probability $1-\rho$.
To address this limitation, Chase et al. [CMY23] proposed a seemingly stronger notion than stability, termed list replicability, which ensures both low population loss and stability of the output hypothesis.
Definition 1.7 (List Replicability).
@@
Similar to stability, we define the notions of global list-replicability and agnostic global list-replicability to describe the uniform case where the learner in Definition 1.7 exists for a fixed constant $L>0$ independent of $\epsilon$.
@@
Surprisingly, Chase, Moran, and Yehudayoff [CMY23] showed that the converse is also true: global stability implies global list-replicability.
@@
For any fixed constant $L>0$, a hypothesis class is $L$-list-replicable if and only if it is $\rho$-stable for all $\rho<\frac{1}{L}$.
@@
Our next theorem extends Theorem 1.8 to show that these new notions coincide with their stability counterparts. Consequently, Theorem 1.4 and Theorem 1.6 completely resolve the questions posed by [CCMY24].
@@
    Given $L:(0,1)\to\mathbb{N}$, if a class $\mathcal{H}$ is $\gamma$-agnostically $L$-list replicable, then it is $\gamma$-agnostically $\rho$-stable for any $\rho:(0,1)\to[0,1]$ satisfying $\rho(\epsilon)<\frac{1}{L(\epsilon)}$ for all $\epsilon\in(0,1)$.
@@
    Given $\rho:(0,1)\to(0,1]$, if a class $\mathcal{H}$ is $\gamma$-agnostically $\rho$-stable, then it is $\gamma$-agnostically $L$-list replicable, for $L(\epsilon)\coloneqq\left\lfloor\frac{1}{\rho(\epsilon/4)}\right\rfloor$.
@@ 1.1 Preliminaries: VC Dimension, Uniform Convergence, and Littlestone Dimension
For the reader’s convenience, we recall the statement of the theorem.
@@
The proof of (i) is straightforward. Given $\varepsilon>0$, let $\delta>0$ be arbitrary and let $n=n(\epsilon,\delta)$ be the sample complexity of a $\gamma$-agnostic $L$-list-replicable learner for $\mathcal{H}$. Let $\mathcal{D}$ be a distribution with population loss at most $\gamma$, and let $h_{1},\ldots,h_{L(\epsilon)}$ be the list of hypotheses satisfying Equation 4. At least one of these hypothesis $h_{i}$ satisfies
@@
Since this statement holds for every $\delta>0$, $\mathcal{H}$ is $\gamma$-agnostically $\rho$-stable for all $\rho(\epsilon)<\frac{1}{L(\epsilon)}$.
@@
Let $n_{0}=n_{0}(\rho,\epsilon)$ be sufficiently large such that the stability property holds, namely, for every $\mathcal{D}$ with population loss at most $\gamma$, there exists $h^{*}:X\to\{0,1\}$ satisfying
@@
By Equations 6 and 7 and the union bound, with probability at least $1-\delta$ , we have
@@
By the equivalence of stability and list-replicability established in Theorem 1.9, Theorem 1.4 (i) is equivalent to the following theorem.
@@
In other words, we sample $\bm{x}$ uniformly at random from $[M]$ and label it according to the hypothesis $\mathbf{1}_{[x\geq m^{*}]}$. We will show that the learner $\mathcal{A}$ cannot satisfy Equations 10 and 11 for this distribution $\mathcal{D}$, resulting in a contradiction.
@@
For every $t\in\{0,\ldots,n\}$, define
@@
    For every pair of distinct functions $g_{1},g_{2}\in G$, there exists $x$ such that $|g_{1}(x)-g_{2}(x)|\geq\delta$.
@@
$\|f_{S}-g\|_{\infty}\leq\delta+\sqrt{\delta}=2\sqrt{\delta}.$
@@
Let $A\subseteq[M/2]$ be the set of $a\in[M/2]$ for which there exists a good sample $S\in\Pi_{t_{0}}$ such that the $c$-th smallest element of $S_{X}$ equals $a$. For every $a\in A$, let $S^{a}$ represent an arbitrary choice of such a good sample. Combining Equations 14 and 17, we have $|A|=\Omega_{n,\epsilon}(M)$.
@@
We show that the algorithm of [BLM20] is already essentially an agnostic $\rho$-stable learner for these classes for some $\rho:(0,1)\rightarrow(0,1)$. For a minor technical reason, in the following lemma, we require that the population loss of the distribution $\mathcal{D}$ is bounded away from $1$. The constant $2/3$ in the statement of the lemma is quite arbitrary and can be replaced by any larger constant strictly less than $1$.
Lemma 4.2 (Stable Agnostic Learning of Classes with Finite Littlestone Dimension).
@@
If $\mathcal{L}_{\mathcal{D}}(\mathcal{H})\geq\frac{2}{3}$, then we have stability, since at least one of the all-$1$ or the all-00 hypothesis has population loss at most $\frac{1}{2}$ which is smaller than $\mathcal{L}_{\mathcal{D}}(\mathcal{H})$. Otherwise, the guarantee of Lemma 4.2 ensures stability. ∎

### v3

@@ 1 Introduction
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\bm{S}))\leq\mathcal{L}_{\mathcal{D}}(\mathcal{H})+\epsilon]\geq 1-\delta\ \text{where }n=n(\epsilon,\delta).$ (1)
@@
Replicability and Global Stability.
Replicability is a fundamental principle of the scientific method. A study is replicable if it consistently yields the same results when repeated with new data drawn from the same distribution or source. In recent years, machine learning has seen a growing need to address the replication crisis [Bal23, Bak16]. Impagliazzo, Lei, Pitassi, and Sorrel [ILPS22] initiated a formal theoretical framework for studying replicability in machine learning. Since their work, a rapidly growing body of research has emerged that introduced various notions of replicability. These works and subsequent research showed that many of these notions of replicability are essentially equivalent. Furthermore, they established deep connections to other foundational concepts in learning theory, such as differential privacy [CMY23, BGH+23, KKMV23, GKM21, CCMY24]. Additionally, a growing body of work has explored replicability in many data analysis and learning settings [ILPS22, BGH+23, KVYZ23, EKK+23, EKM+23, EHKS23, KKL+24, KKMV23].
In this paper, we focus on the notion of replicability where the learning algorithm is expected to often produce the same predictor when applied to two independent and identically distributed inputs. This concept was first introduced under the term global stability in [BLM20] and has since been refined and explored in subsequent works [GKM21, KKMV23, CMY23, CCMY24]. We start by defining global stability.
Definition 1.1 ($\rho$-Global Stability, [CCMY24]).
Given a function $\rho:(0,1)\rightarrow(0,1)$, a learning rule $\bm{\mathcal{A}}$ is a $\rho$-global stable learner for a hypothesis class $\mathcal{H}$ if the following holds. For every $\epsilon>0$, there exists $n=n(\epsilon)$ such that for every realizable distribution $\mathcal{D}$, there exists a hypothesis $h$ satisfying
@@
Similarly, we call $\bm{\mathcal{A}}$ a $\rho$-global stable agnostic learner for $\mathcal{H}$, if there exists $n=n(\epsilon)$ such that for every distribution $\mathcal{D}$ on $X\times\{0,1\}$, there exists a hypothesis $h\in\{0,1\}^{X}$ that satisfies (2) and
@@
To simplify terminology, we use the term $\rho$-global stable to describe a hypothesis class $\mathcal{H}$ with a $\rho$-global stable learner. Likewise, we call $\mathcal{H}$ agnostically $\rho$-global stable if it has a $\rho$-global stable agnostic learner.
Definition 1.2 (Global Stability, [BLM20]).
We say that a hypothesis class $\mathcal{H}$ is globally stable if it is $\rho$-global stable for a fixed constant $\rho\in(0,1)$. Similarly, a hypothesis class $\mathcal{H}$ is agnostically globally stable if it is agnostically $\rho$-stable for such a constant.
@@
Bun, Livni, and Moran [BLM20] showed that in the realizable setting, global stability is fully characterized by bounded Littlestone dimension (Definition 1.12). This result, combined with the seminal works of Littlestone [Lit88] and Alon et al. [ABL+22], shows that global stability is equivalent to online learnability and approximately private learnability, as well as some other notions of replicability [KKMV23, GKM21, BGH+23].
@@
This striking result shows that agnostic global stability is far more restrictive than its realizable counterpart. Since finite classes are trivially global stable, Theorem 1.3 shows that agnostic global stability is too restrictive to lead to interesting learnability phenomena. To remedy this, Chase et al. [CCMY24] introduced two relaxations of agnostic global stability and proposed a study of which hypothesis classes can be learned under these relaxed notions of stability.
@@
The first suggested relaxation, coincides with our definition of $\rho$-global stability in Definition 1.1. A hypothesis class $\mathcal{H}$ is called excess-error dependent stable if it is agnostically $\rho$-global stable for some $\rho:(0,1)\to(0,1)$. Here, the excess-error refers to the parameter $\epsilon$ in Eq. 3.
Our main theorem provides a complete characterization of such classes. We show that a hypothesis class is agnostically $\rho$-global stable learnable for some $\rho$ if and only if it has a bounded Littlestone dimension. We denote the Littlestone dimension of $\mathcal{H}$ as $\operatorname{Ldim}(\mathcal{H})$.
@@
    If $\operatorname{Ldim}(\mathcal{H})=\infty$, then $\mathcal{H}$ is not $\rho$-global stable for any $\rho:(0,1)\to(0,1)$.
@@
    If $\operatorname{Ldim}(\mathcal{H})<\infty$, then $\mathcal{H}$ is agnostically $\rho$-global stable for some $\rho:(0,1)\to(0,1)$.
Note that Theorem 1.4 (i) states that if $\operatorname{Ldim}(\mathcal{H})=\infty$, then even in the realizable case, we cannot achieve $\rho$-global stability for any $\rho:(0,1)\to(0,1)$. This strengthens the result of Bun, Livni, and Moran [BLM20], which only overrules $\rho$-global stability when $\rho>0$ is a fixed constant.
Shortly after a draft of this paper was posted online, Hopkins and Moran [HM25] communicated to us that in an independent work, they have proved an equivalent statement to Theorem 4 by utilizing the known relation between stability and differential privacy. They use a $\rho$-global stable learner to achieve weak DP learning, which in turn is boosted to a strong DP learner. It is well known that strong DP learning is achievable if and only if the Littlestone dimension is finite [ABL+22]. In contrast, our proof is direct and relies solely on notions of stability and list-replicability.
Combined with the work of Alon et al. [ABL+22], Theorem 1.4 implies that agnostic $\rho$-global stability is equivalent to global stability, as well as to approximate private learnability and online learnability.
@@
Definition 1.5 (Class-error Dependent Stability, [CCMY24]).
@@
To prove our theorem, we show that agnostic global stability reduces to $\gamma$-agnostic global stability, for any arbitrary $\gamma>0$.
@@
In learning theory, global stability is more useful when paired with a guarantee that the learner typically outputs a hypothesis with a low population loss. The definition of global stability (Definition 1.2) only requires that the learner outputs a low-error hypothesis with some probability $\rho>0$, and the learner can output hypotheses with large population loss with probability $1-\rho$. However, it is known that global stability implies bounded VC dimension [ABL+22], and assuming bounded VC dimension, this can be easily remedied. The learner can estimate the population loss of its output by comparing it to that of the hypothesis produced by the empirical risk minimization (ERM) rule. If the population loss is unsatisfactory, the learner can fall back on the ERM hypothesis instead.
Next, we introduce a seemingly stronger notion of replicability, originally proposed by Chase et al. [CMY23], known as list replicability. This notion strengthens the standard guarantee by requiring that, with high probability, the output hypothesis lies within a small list of hypotheses, each with low population loss.
Definition 1.7 (List Replicability, [CMY23]).
@@
Similar to stability, we define the notions of global list-replicability and agnostic global list-replicability to describe the uniform case where the learner in Definition 1.7 exists for a fixed constant $L>0$ independent of $\epsilon$ [^1] In the literature, what we refer to as global list-replicability is simply called list-replicability.
@@
Chase, Moran, and Yehudayoff [CMY23] showed that the converse is also true: global stability implies global list-replicability.
@@
For any fixed constant $L>0$, a hypothesis class is $L$-list-replicable if and only if it is $\rho$-global stable for all $\rho<\frac{1}{L}$.
@@
Our next theorem extends Theorem 1.8 to show that these new notions coincide with their stability counterparts. Consequently, Theorem 1.4 and Theorem 1.6 completely resolve the questions posed in [CCMY24].
@@
    Given $L:(0,1)\to\mathbb{N}$, if a class $\mathcal{H}$ is $\gamma$-agnostically $L$-list replicable, then it is $\gamma$-agnostically $\rho$-global stable for any $\rho:(0,1)\to[0,1]$ satisfying $\rho(\epsilon)<\frac{1}{L(\epsilon)}$ for all $\epsilon\in(0,1)$.
@@
    Given $\rho:(0,1)\to(0,1]$, if a class $\mathcal{H}$ is $\gamma$-agnostically $\rho$-global stable, then it is $\gamma$-agnostically $L$-list replicable, for $L(\epsilon)\coloneqq\left\lfloor\frac{1}{\rho(\epsilon/4)}\right\rfloor$.
@@ 1.1 Preliminaries: VC Dimension, Uniform Convergence, and Littlestone Dimension
For the reader’s convenience, we recall the statement of the theorem and provide a summary of the proof.
@@
To prove (ii), we construct an agnostic $L$-list replicable learner by running the $\rho$-global stable learning algorithm multiple times. We return any output hypothesis whose empirical loss is close to that of the best output hypothesis, and whose empirical frequency is not much smaller than $\rho$. This guarantees that we typically output a hypothesis with low population loss and high likelihood of being an output of the globally stable learner. The latter ensures that our output is typically confined to a small list.
@@
The proof of (i) is straightforward. Given $\varepsilon>0$, let $\delta>0$ be arbitrary and let $n=n(\epsilon,\delta)$ be the sample complexity of a $\gamma$-agnostic $L$-list-replicable learner for $\mathcal{H}$. Let $\mathcal{D}$ be a distribution with population loss at most $\gamma$, and let $h_{1},\ldots,h_{L(\epsilon)}$ be the list of hypotheses satisfying Equation 5. At least one of these hypotheses $h_{i}$ satisfies
@@
Since this statement holds for every $\delta>0$, $\mathcal{H}$ is $\gamma$-agnostically $\rho$-global stable for all $\rho(\epsilon)<\frac{1}{L(\epsilon)}$.
@@
Let $n_{0}=n_{0}(\rho,\epsilon)$ be sufficiently large such that the global stability property holds, namely, for every $\mathcal{D}$ with population loss at most $\gamma$, there exists $h^{*}:X\to\{0,1\}$ satisfying
@@
By Equations 7 and 8 and the union bound, with probability at least $1-\delta$ , we have
@@
By the equivalence of global stability and list-replicability established in Theorem 1.9, Theorem 1.4 (i) is equivalent to the following theorem.
@@
Let us give an overview of the proof of Theorem 3.1. By Proposition 3.3, it is sufficient to prove that hypothesis classes of infinite threshold dimension are not $L$-list replicable. Assume towards contradiction that there exists an $L$-list-replicable learner for a hypothesis class of infinite threshold dimension.
Similarly to Alon et al. [ALMM19], we use a hypergraph Ramsey argument to restrict the learning problem to an arbitrarily large subset $X$ of the domain, on which the learner’s prediction is essentially determined by the ordered sign pattern of the sample. In particular, the Ramsey argument ensures that to label an element $x\in X$, the learner essentially looks at the ordered sign pattern of the labelled points in its given sample, and also at the “order” of $x$ with respect to these points.
We consider the distribution that is uniform over the threshold about the median of $X$. Since the learner is a PAC learner, we show that the probability that its output classifies a point $x$ as a $1$ ranges from close to 00 to close to $1$, as the order of $x$ ranges from the smallest to the largest element compared to the points in the samples. Consequently, we detect a probability jump from some order to the next.
The most significant difference between our proof and the proof of [ALMM19] lies in handling this probability jump. [ALMM19] exploits the jump to create a “privacy leak,” whereas, without a privacy guarantee, we take a different approach based on an “approximate rank” argument. More specifically, we use our boosting lemma (Lemma 3.5) to show that for many samples, the function corresponding to the output probabilities can be well-approximated (in the $L_{\infty}$ norm) by a convex combination of a fixed short list of hypotheses. We then apply a “volume-based” argument to derive a contradiction by finding two samples that are well-approximated by the same convex combination, but are supposed to label a point $x\in X$ differently according to the aforementioned probability jump.
@@
In other words, we sample $\bm{x}$ uniformly at random from $[M]$ and label it according to the hypothesis $\mathbf{1}_{[x\geq m^{*}]}$. We will show that the learner $\mathcal{A}$ cannot satisfy Equations 11 and 12 for this distribution $\mathcal{D}$, resulting in a contradiction.
@@
For every $t\in[\frac{n}{4},\frac{3n}{4}]$, define
@@
    For every pair of distinct functions $g_{1},g_{2}\in G$, there exists $x\in X^{\prime}$ such that $|g_{1}(x)-g_{2}(x)|\geq\delta$.
@@
$\|f_{S}-g\|_{\infty}\leq\delta+\sqrt{\delta}\leq 2\sqrt{\delta}.$
@@
Let $A\subseteq[M/2]$ be the set of $a\in[M/2]$ for which there exists a good sample $S\in\Pi_{t_{0}}$ such that the $c$-th smallest element of $S_{X}$ equals $a$. For every $a\in A$, let $S^{a}$ represent an arbitrary choice of such a good sample. Combining Equations 15 and 18, we have $|A|=\Omega_{n,\epsilon}(M)$.
@@
We show that the algorithm of [BLM20] is already essentially an agnostic $\rho$-global stable learner for these classes for some $\rho:(0,1)\rightarrow(0,1)$. For a minor technical reason, in the following lemma, we require that the population loss of the distribution $\mathcal{D}$ is bounded away from $1$. The constant $2/3$ in the statement of the lemma is quite arbitrary and can be replaced by any larger constant strictly less than $1$.
Lemma 4.2 (Agnostic $\rho$-Global Stable Learning of Classes with Finite Littlestone Dimension).
@@
If $\mathcal{L}_{\mathcal{D}}(\mathcal{H})\geq\frac{2}{3}$, then we have global stability, since at least one of the all-$1$ or the all-00 hypothesis has population loss at most $\frac{1}{2}$ which is smaller than $\mathcal{L}_{\mathcal{D}}(\mathcal{H})$. Otherwise, the guarantee of Lemma 4.2 ensures stability. ∎